# CRPS-Based Targeted Sequential Design with Application in Chemical Space

### v1

@@ 5.2 Original dataset
For the original dataset, we only consider the box and whiskers plot for comparison (Fig. 7), with focusing solely on (a) $\overline{\mathrm{CRPS}}_{\nu_{1}}$ and (b) $\overline{\mathrm{CRPS}}_{\nu_{2}}$ (as the ground truth is not available, see Section 4.4). The pointwise $\mathrm{CRPS}_{\nu_{1}}$ also performs best for $\overline{\mathrm{CRPS}}_{\nu_{1}}$ in the original dataset and reduces the median value of the random strategy by 25 %. For $\overline{\mathrm{CRPS}}_{\nu_{2}}$, as in the synthetic case, $\mathrm{ICRPS}_{\nu_{2}}$ and TIMSE lead to similar results, with a slightly lower median for $\mathrm{ICRPS}_{\nu_{2}}$, which reduces the value of the random strategy by 15 %.
Figure 7: Comparison of the data acquisition methods using the original dataset and $M=100$ validation molecules, 60 repetitions, $n=30$ and $\widetilde{n}=25$. The box and whiskers plots depict the median, quartiles, two whiskers and all ’outlying’ points. The whiskers extend from the quartiles to the smallest / highest value at most 1.5 $\times$ the interquartile range away from the quartile, beyond which points are considered outliers.
@@ 6 Discussion and Conclusions
The test case using the synthetic version of the Photoswitch dataset highlights the following strengths and weaknesses of the new selection criteria: In terms of classification, $\mathrm{CRPS}{\nu_{1}}$ and $\mathrm{ICRPS}{\nu_{1}}$ are more precise, while $\mathrm{CRPS}{\nu_{2}}$ and $\mathrm{ICRPS}{\nu_{2}}$ are more sensitive (Fig. 3 c and f). For precision, $\mathrm{CRPS}{\nu_{1}}$ performs best, while $\mathrm{ICRPS}{\nu_{1}}$ remains precise but also outperforms the random approach in sensitivity. Regarding $\mathrm{CRPS}{\nu_{2}}$, the pointwise criterion is both more precise and sensitive. Regarding the prediction for the points identified as part of the excursion set, $\mathrm{CRPS}_{\nu_{1}}$ is highly accurate, both when considering the mean alone ($\mathrm{RMSE}_{p}$, Fig. 3b) and the predictive distribution ($\overline{\mathrm{CRPS}}_{\nu_{1}}$, Fig. 3a). But again, as for precision, $\mathrm{ICRPS}_{\nu_{1}}$ shows a similar performance and is at the same time more sensitive. Considering the points actually lying in the excursion set, $\mathrm{ICRPS}_{\nu_{2}}$ shows the strongest performance ($\overline{\mathrm{CRPS}}_{\nu_{2}}$ and $\mathrm{RMSE}_{\Gamma}$, Fig. 3d+e). Focusing on sensitivity only, $\mathrm{CRPS}{\nu_{2}}$ performs better, but $\mathrm{ICRPS}{\nu_{2}}$ does a better job when also considering the full predictive distribution.
@@
When comparing the new criteria to state-of-the-art methods (Fig. 6), we observe that the performance of $\mathrm{CRPS}{\nu_{2}}$ and $\mathrm{ICRPS}{\nu_{2}}$ is similar to the existing TMSE and TIMSE criteria, even in terms of $\overline{\mathrm{CRPS}}_{\nu_{2}}$. This is expected, as all these methods rely on a Gaussian PDF around the threshold for targeted acquisition. On the other hand, $\mathrm{CRPS}{\nu_{1}}$ and $\mathrm{ICRPS}{\nu_{1}}$ outperform the other criteria in terms of $\overline{\mathrm{CRPS}}_{\nu_{1}}$, $\mathrm{RMSE}_{p}$, and precision (Fig. 3a-c). Entropy performed badly across all criteria, with both entropy and IBV showing notably poor performance in $\mathrm{RMSE}_{\Gamma}$ (Fig. 6e). In the real test case, where no ground truth was available (Fig. 7), we relied on $\overline{\mathrm{CRPS}}_{\nu_{1}}$ and $\overline{\mathrm{CRPS}}_{\nu_{2}}$, observing results similar to those in the synthetic test case, which helped confirm the conclusions.
@@
There are several directions in which the presented work can be extended and deepened. Other acquisition criteria, related to expected scoring rules, can also be explored. For instance, the entropy-based criterion (Cole et al.,, 2023) is equivalent to the expected log score for a binary predictor. Similarly, the integrand of the IBV is the Bernoulli variance, which corresponds to the expectation of the Brier Score (Brier,, 1950; Murphy,, 1973). Future work could explore further alternatives to the CRPS, including multivariate scoring rules (e.g., Allen et al., 2023b, ). Within the threshold-weighted CRPS, alternative weighting measures could be explored, such as the corresponding CDF of the Gaussian $\nu_{2}$ (Allen,, 2024). This would result in a criterion that is less similar to TMSE and TIMSE. As for the criteria based on $\mathrm{CRPS}_{\nu_{1}}$, we could adjust the threshold in the weighting measure to incorporate the region around the threshold, thereby enhancing the method’s sensitivity. Next, k-step look-ahead strategies could be explored as an alternative to myopic ones. Furthermore, the initial design is currently random, what could be leveraged by employing a space-filling approach.
Acknowledgements
Declaration of Interest
The authors report there are no competing interests to declare.

### v2

@@ 5.2 Original dataset
For the original dataset, we only consider the box and whiskers plot for comparison (Fig. 7), with focusing solely on (a) $\overline{\mathrm{CRPS}}_{\gamma_{1}}$, (b) $\overline{\mathrm{CRPS}}_{\gamma_{2}}$ and (c) $\overline{\mathrm{CRPS}}$ (as the ground truth is not available, see Section 4.4). The pointwise $\mathrm{CRPS}_{\gamma_{1}}$ also performs best for $\overline{\mathrm{CRPS}}_{\gamma_{1}}$ in the original dataset and reduces the median value of the random strategy by 30 %. For $\overline{\mathrm{CRPS}}_{\gamma_{2}}$, as in the synthetic case, $\mathrm{ICRPS}_{\gamma_{2}}$ and TIMSE lead to similar results, reducing the value of the random strategy by 15 %. As in the synthetic test case, the lowest value for $\overline{\mathrm{CRPS}}$ is also obtained by $\mathrm{ICRPS}_{\gamma_{2}}$ and TIMSE.
Figure 7: Comparison of the data acquisition methods using the original dataset and $M=100$ validation molecules, 100 repetitions, $n=30$ and $\widetilde{n}=25$. The box and whiskers plots depict the median, quartiles, two whiskers and all ’outlying’ points. The whiskers extend from the quartiles to the smallest / highest value at most 1.5 $\times$ the interquartile range away from the quartile, beyond which points are considered outliers.
@@ 6 Discussion and Conclusions
The test case using the synthetic version of the Photoswitch dataset highlights the following strengths and weaknesses of the new selection criteria: In terms of classification, $\mathrm{CRPS}_{\gamma_{1}}$ is the most precise, while $\mathrm{CRPS}_{\gamma_{2}}$ and $\mathrm{ICRPS}_{\gamma_{2}}$ are more sensitive (Fig. 3 c and f). Regarding the prediction for the points identified as part of the excursion set, $\mathrm{CRPS}_{\gamma_{1}}$ is highly accurate, both when considering the mean alone ($\mathrm{RMSE}_{p}$, Fig. 3b) and the predictive distribution ($\overline{\mathrm{CRPS}}_{\gamma_{1}}$, Fig. 3a). However, $\mathrm{ICRPS}_{\gamma_{1}}$ shows a comparable performance and is at the same time more sensitive. Considering the points actually lying in the excursion set, $\mathrm{ICRPS}_{\gamma_{2}}$ shows overall the strongest performance ($\overline{\mathrm{CRPS}}_{\gamma_{2}}$$\mathrm{RMSE}_{\Gamma}$ and sensitivity, Fig. 3d-f).
@@
When comparing the new criteria to state-of-the-art methods (Fig. 6), we observe that the performance of $\mathrm{CRPS}_{\gamma_{2}}$ and $\mathrm{ICRPS}_{\gamma_{2}}$ is similar to the existing TMSE and TIMSE criteria, even in terms of $\overline{\mathrm{CRPS}}_{\gamma_{2}}$. This is expected, as all these methods rely on a Gaussian PDF around the threshold for targeted acquisition. On the other hand, $\mathrm{CRPS}_{\gamma_{1}}$ and $\mathrm{ICRPS}_{\gamma_{1}}$ clearly outperform the other criteria in terms of $\overline{\mathrm{CRPS}}_{\gamma_{1}}$ and $\mathrm{RMSE}_{p}$ (Fig. 3a+b). In the real test case, where no ground truth was available (Fig. 7), we relied on $\overline{\mathrm{CRPS}}_{\gamma_{1}}$, $\overline{\mathrm{CRPS}}_{\gamma_{2}}$ and $\overline{\mathrm{CRPS}}$, observing results similar to those in the synthetic test case, which helped confirm the conclusions.
@@
There are several directions in which the presented work can be extended and deepened. Other acquisition criteria, related to expected scoring rules, can also be explored. For instance, the entropy-based criterion (Cole et al.,, 2023) is equivalent to the expected log score for a binary predictor. Similarly, the integrand of the IBV is the Bernoulli variance, which corresponds to the expectation of the Brier Score (Brier,, 1950; Murphy,, 1973). Future work could explore further alternatives to the CRPS, including multivariate scoring rules (e.g., Allen et al., 2023b, ). Within the threshold-weighted CRPS, alternative weighting measures could be explored, such as the corresponding CDF of the Gaussian $\gamma_{2}$ (Allen,, 2024). This would result in a criterion that is less similar to TMSE and TIMSE. As for the criteria based on $\mathrm{CRPS}_{\gamma_{1}}$, we could adjust the threshold in the weighting measure to incorporate the region around the threshold, thereby enhancing the method’s sensitivity. Next, k-step look-ahead strategies could be explored as an alternative to myopic ones. Furthermore, the initial design is currently random, what could be leveraged by employing a space-filling approach.
Supplementary Material
PDF containing (1) more on Gaussian Process modeling, (2) a kernel comparison with the (original) Photoswitch dataset, (3) proofs of the analytical expressions for the threshold-weighted CRPS and (4) a proof of consistency for the SUR strategy.